Minimize SLSTR Calibration

Signals from the VISCAL source for the VIS channels show oscillations due to the build up of ice on the optical path within the FPA.

Decontamination is carried out periodically, in order to warm up the FPA and remove the ice. The last decontamination cycle was successfully performed following the anomaly on 14th February 2017.

Following an anomaly on 16th March 2017 and the subsequent cooldown, the VISCAL signals have returned to their pre-anomaly levels and the oscillations in signal have continued from where they left off.

 

VIS

VISCAL signal trend for VIS channels (nadir view)

 

SWIR

VISCAL signal trend for SWIR channels (nadir view)